Ask
your own question, for FREE!
Mathematics
16 Online
Sue chooses a dress and necklace to wear when she performs at a violin recital. The probability that Sue selects a black dress is 40%. The probability Sue selects a gold necklace is 30%. What is the probability that Sue chooses a black dress and a gold necklace? A. 12% B. 35% C. 50% D. 70%
Still Need Help?
Join the QuestionCove community and study together with friends!
Hello, welcome to QuestionCove cx You have to multiple the percents (I out them into fraction form) 4/10*3/10=12/100 which is 12% 12% chance of getting a black dress and a gold necklace
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
Kimmee1001:
Who or what is ultimately responsible for Romeo and Juliet's deaths?
EdwinJsHispanic:
is it normal to have high expectations, like expecting your partner to have time
gelphielvr:
Advice for cramming 70 vocab terms each week? Taking med term this year and we have to memorize about 70+ vocab terms per week and im having trouble studyin
EdwinJsHispanic:
I lowk need help deciding what I should write/type on this essay about why teache
GRIZZLEY19:
Let me know what y'all think about the story Image waking up and seeing a family in the dark.
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
12 hours ago
1 Reply
0 Medals
1 day ago
4 Replies
1 Medal
1 day ago
4 Replies
0 Medals
17 hours ago
3 Replies
0 Medals
1 day ago
22 Replies
6 Medals
1 week ago
1 Reply
0 Medals